If the kinetic energy of an electron in the ground state hydrogen atom is 2.2 x 10-18 J, calculate its momentum.


1
Expert's answer
2020-07-02T17:14:18-0400
"p=\sqrt{2m_eK}\\\implies p=2.0\times 10^{-24}kg\cdot m/s"